# Accuracy and Political Bias of News Source Credibility Ratings by Large Language Models

### v2

@@ Discussion
Our study only scratches the surface when it comes to the impact of LLMs as information curators. Future studies should explore how LLMs handle information from different sources in more realistic scenarios, how other information selection criteria, such as relevance and freshness (Sun et al. 2023), affect the outcomes, and how to mitigate emerging manipulation techniques (Pfrommer et al. 2024). At the same time, more studies on how humans interact with AI-augmented systems (Sharma, Liao, and Xiao 2024; DeVerna et al. 2024) and how information processed by AI influences humans are needed (Jakesch et al. 2023; Simchon, Edwards, and Lewandowsky 2024).
Appendix A Appendix
Complete Prompts
@@ Appendix A Appendix
System prompt You are an assistant to determine the credibility of websites.
@@
Task instruction Rate the website’s credibility: {domain}, on a scale between 0 and 1, where 0 means very low credibility and 1 means very high credibility. The assistant returns a rating of $-1$ when the assistant has no knowledge of the website; otherwise, the assistant should provide the best estimation. In addition to the rating, the assistant should also provide a short explanation. The output should be in JSON format and follow the following schema. Only return the JSON object, nothing else. {‘url’: ‘example.com’, ‘rating’: 0.5, ‘explanation’: ‘The example website is known to post credible content.’}
@@
When testing the impact of political identities, we modify the system prompt:
Political identity assignment You are an assistant to determine the credibility of websites. You identify as {identity} on the US political spectrum.
In the experiments, {identity} is replaced with the one of “a Democrat,” “an Independent,” and “a Republican.” The task instruction remains the same.
Response Examples
@@
Table 1: Ratings of reuters.com by different LLMs.
@@
Table 2: Ratings of aldf.com by different LLMs. Five models believe the website belongs to the Animal Legal Defense Fund (aldf.org).
@@
Gemini 1.5 Flash -1 I have no information about this website.

### v3

@@ 5.4. Future Studies
Other important directions include exploring how LLMs handle information from different sources in more realistic scenarios, how other information selection criteria, such as relevance and freshness (Sun et al., 2023), affect the outcomes, and how to mitigate emerging manipulation techniques (Pfrommer et al., 2024). At the same time, more studies on how humans interact with AI-augmented systems (Sharma et al., 2024; DeVerna et al., 2024) and how information processed by AI influences humans are needed (Jakesch et al., 2023; Simchon et al., 2024).
Appendix A Complete Prompts
@@
System prompt:
You are an assistant to determine the credibility of websites.
@@
Task instruction:
Rate the website’s credibility: {domain}, on a scale between 0 and 1, where 0 means very low credibility and 1 means very high credibility. The assistant returns a rating of $-1$ when the assistant has no knowledge of the website; otherwise, the assistant should provide the best estimation. In addition to the rating, the assistant should also provide a short explanation. The output should be in JSON format and follow the following schema. Only return the JSON object, nothing else. {‘url’: ‘example.com’, ‘rating’: 0.5, ‘explanation’: ‘The example website is known to post credible content.’}
@@
When testing the impact of partisan roles, we modify the system prompt:
Partisan role assignment:
You are an assistant to determine the credibility of websites. You identify as {role} on the US political spectrum.
In the experiments, {role} is replaced with the one of “a Democrat,” “an Independent,” or “a Republican.” The task instruction remains the same.
Appendix B Response Examples
@@
Table 1. Ratings of reuters.com by different LLMs.
@@
o1-mini 0.9 Reuters is a well-established news organization known for reliable and unbiased reporting.
@@
Table 2. Ratings of aldf.com by different LLMs. Five models believe the website belongs to the Animal Legal Defense Fund (aldf.org).
@@
o1-mini 0.7 The website aldf.com is associated with the American Lebanese Democratic Federation, a recognized organization, indicating a moderate to high level of credibility.
Gemini 1.5 Flash -1 I have no information about this website.
@@
Appendix C Additional Results
Figure 9. Distribution of domain credibility ratings from different LLMs and human experts.
[…]
Distribution of domain credibility ratings from different LLMs and human experts.
In Figure 9, we show the distribution of domain credibility ratings from different LLMs and human experts.